How much steam at 100°C must be added to 163g of water at 250C contained in a 30 g glass vessel to produce water at 77°C?

1
Expert's answer
2021-11-24T11:59:59-0500

Find the heat required:


"m(r+c_w(t_s-t_f))=(m_wc_w+m_gc_g)(t_f-t_w),\\\space\\
m=\frac{(m_wc_w+m_gc_g)(t_f-t_w)}{r+c_w(t_s-t_f)}=0.016\text{ kg}."